Summary for Eve shift, Nov. 28, 1999

Angela Bellavance (bellavan@rice.edu)
Mon, 29 Nov 1999 00:02:10 -0600 (CST)

Shift Summary for eve shift, 28/Nov/99

On Shift: A. Tripathi, A. Bellavance

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0.8

Accounting of lost time: Recovery from power outages.

========================================================================
Runs ended this shift:

15018 - LASER_SCAN

Tapes brought to FCC this shift: none, but there are some waiting to be
taken tomorrow, day shift.
(including tapes for online split)

========================================================================
Accesses during shift: open access

========================================================================
Overview of shift:
Flurry of work by a lot of people to get the detector functioning again
properly after the power outages of this morning. Estimate for beam
return is currently 7am.

========================================================================
Description of Shift:

time comment
------------------------------------------------------------------------
15:40 Onsplit restarted, CsI RF system restarted, pipeline inited.
16:01 Epicure restarted, ZCal moved to out position. Magnet
re-setting will have to wait until interlocks are
re-established.
17:12 Pulled VV' crate 4-51 out of the readout. This will need to be
going again to take data. Seems like the TDC in slot 9 has
failed.
17:15 Two new bottles of Ar-Eth now online, west side.
18:15 Ashkan futzes with the cabling for DC4X_H_R. See logbook for
details, will need follow-up work.
18:45 Ion gauge alarm goes off. Doug J. called in. Blower pump 1 had
tripped off. No cause found, blower reset, seems fine.
20:23 Leo: "Time for another fireball... Put that in the shift
summary..."
22:00 Gas checklist done. Flowmeter in gas shed for DC2 is way too
low. Dave N. paged. Flowmeter downstairs looks fine, so rely
on that one for now.
23:00 Crew shows up to fix interlocks; done. Call them back to S&S.

========================================================================
Items to Follow up on Next Shift

- need search and secure.
- open valve from NM3 to NM4. (Doug J.)
- work on getting fruit back into readout. (Jason L.)
- reset all Epicure magnets (including turning on NM4AN); do muon run
after flipping polarity to NEG.
- turn in tapes to Feynman (Monday day shift).

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke